# A Survey on the Applications of Zero-Knowledge Proofs

### v1

@@ Abstract
Zero-knowledge proofs (ZKPs) represent a revolutionary advance in computational integrity and privacy technology, enabling the secure and private exchange of information without revealing underlying private data. ZKPs have unique advantages in terms of universality and minimal security assumptions when compared to other privacy-sensitive computational methods for distributed systems, such as homomorphic encryption and secure multiparty computation. Their application spans multiple domains, from enhancing privacy in blockchain to facilitating confidential verification of computational tasks. This survey starts with a high-level overview of the technical workings of ZKPs with a focus on an increasingly relevant subset of ZKPs called zk-SNARKS. While there have been prior surveys on the algorithmic and theoretical aspects of ZKPs, our work is distinguished by providing a broader view of practical aspects and describing many recently-developed use cases of ZKPs across various domains. These application domains span blockchain privacy, scaling, storage, and interoperability, as well as non-blockchain applications like voting, authentication, timelocks, and machine learning. Aimed at both practitioners and researchers, the survey also covers foundational components and infrastructure such as zero-knowledge virtual machines (zkVM), domain-specific languages (DSLs), supporting libraries, frameworks, and protocols. We conclude with a discussion on future directions, positioning ZKPs as pivotal in the advancement of cryptographic practices and digital privacy across many applications.
I Introduction
@@
Zero-knowledge proofs (ZKPs) are a set of cryptographic methods allowing one party to prove the validity of a claim to another without disclosing any of the claim’s underlying details. The seminal work by Goldwasser, Micali, and Rackoff laid the groundwork for ZKPs in the 1980s [1]. It introduced the principle of knowledge complexity as a metric for quantifying the information transferred from the prover to the verifier. Subsequently, Goldreich and others [2, 3] published several works until the 1990s that expanded the scope of ZKPs to a broader range of computational problems. They demonstrated their applicability to NP-complete problems under specific cryptographic assumptions. Another significant advancement in the development of ZKP’s came with the introduction of succinctness in [4, 5] to ZKP’s following the design paradigm of Kilian’s seminal 1992 paper [6]. The succinct and non-interactive properties of zk-SNARKs greatly enhance their practical applicability and versatility in blockchains and non-blockchain systems. The foundational principles of ZKPs offer innovative solutions to digital systems that prioritize both security and privacy [7].
@@
The advent of ZKPs has represented a significant leap forward in the convergence of privacy and verifiability, which has captivated significant interest from both the cryptographic scholarly community and the industry. It introduced a unique approach in the field of cryptography, differentiating itself from other privacy-sensitive computational innovations for distributed systems, such as homomorphic encryption and secure multiparty computation. While these methods are also being actively developed and advanced, each serves a specific purpose in information verification and privacy preservation. Homomorphic encryption enables computations on encrypted data without needing to decrypt, thereby preserving confidentiality while allowing the derivation of useful insights [8]. Secure multiparty computation enables trustless collaborations and allows parties to jointly compute a function over their inputs while keeping those inputs private [9]. From Table I, we observe ZKPs offer several advantages, such as universality and minimal security assumptions over other techniques by proving the truth of a statement without disclosing any other information.
Cryptographic Method Universality Security Assumptions Computational Complexity
Zero-Knowledge Proofs High: Versatile across various protocols. Supports a broad spectrum of functions from authentication to smart contracts. Minimal: Relies on the complexity of creating a proof without exposing any of the underlying data. Assumes the security of cryptographic primitives. High: The prover and verifier may perform significant computational work, which might include prover creating a polynomial time computation representing the original problem and generating a proof of this computation
Fully Homomorphic Encryption High: Best suited for secure processing of data. Allows for computation on encrypted data, making it ideal for cloud computing environments. Scheme Dependent: Security is contingent on the difficulty of algebraic challenges. Very High: FHE operations are computationally intensive because they involve complex algebraic operations on encrypted data. The overhead associated with maintaining homomorphism over computations makes these operations costly.
Secure Multiparty Computation Medium: Geared towards collaborative tasks in data analysis. Permits multiple parties to compute over data without revealing individual inputs. Varied: Security often needs an honest majority within the group of participants, assuming computational hardness and sometimes the use of random oracles. High: Depends on protocol complexity and the number of participants. Each participant in an SMC protocol must perform computations on their part of the data as well as on data received from others.
TABLE I: Comparison of Zero-Knowledge Proofs with other cryptographic methods
@@
In modern digital systems, a trade-off exists between openness and privacy. Blockchains, for example, prioritize transparency to ensure trust and prevent fraud, with every transaction openly verifiable. However, this transparency can compromise privacy [10]. Despite the pseudonymous nature of blockchain transactions, advanced analytics can de-anonymize users by correlating on-chain and off-chain data. Such exposure can reveal a user’s entire transaction history, leading to privacy breaches and potential targeted threats. As digital infrastructures become increasingly complex, striking the right balance between transparency for security and preserving user privacy becomes a vital research challenge. In this context, ZKPs emerge as a robust solution addressing the challenges posed by such trade-offs in digital systems.
@@
ZKPs enable users to verify private data such as bank balances or credit scores without revealing specifics. Furthermore, they can ensure anonymity in authorization processes by allowing access to restricted areas or proving regional origin without disclosing detailed credentials. In the financial domain, ZKPs can allow for identity-free payments and tax submissions without revealing exact earnings. Additionally, they facilitate trustless outsourcing by letting organizations validate results without redoing the entire operation in the computational field. Furthermore, ZKPs can modify blockchain operations, shifting from collective to singular computation with network-wide verification. Overall, ZKPs exemplify the merging of verification with privacy across diverse applications [11].
In exploring the realm of ZKPs, it is fundamental to recognize verifiable computation as the cornerstone upon which the practical applications of ZKPs are built. Verifiable computation, at its core, is the ability to prove that an external computation was performed correctly without revealing the inputs or the computation process. This foundational attribute of ZKPs serves as a gateway to their two practical value propositions: succinctness and privacy. In the remainder of this survey, we will see how each of these two values contributes to the practical applications of ZKPs.
@@
Succinctness in ZKPs allows for the quick verification of the correctness of a computation without the extensive resources typically required for direct computation execution. This is crucial in contexts where computational resources are limited or expensive, such as in blockchain networks. For instance, ZKPs enable the validation of transaction blocks or smart contract executions without necessitating each node to replay the entire sequence of transactions or computations. This aspect dramatically reduces the computational load on the network, enhancing scalability and throughput.
Privacy, the second major value proposition, emerges from the intrinsic nature of ZKPs to prove the correctness of information without revealing the information itself. This characteristic is particularly transformative in scenarios where sensitive or confidential data is involved. This capability opens the door to a wide range of applications, from private voting systems to confidential financial transactions, where the integrity of the process is maintained without compromising the privacy of the individuals or entities involved.
@@
In the ZKP literature, there exist two key survey articles relevant to our work. The first survey by Morais et al. [12] dives deep into Zero-Knowledge Range Proofs (ZKRP) with a particular emphasis on the algorithmic details required for the implementation of ZKRPs. Furthermore, it defines the fundamental theoretical background behind ZKRPs and describes possible generic and blockchain-specific use cases of ZKRPs. In contrast, our survey covers a broader range of application categories and use cases. Further, we provide an overview of the underlying theoretical components, tools, and infrastructure needed for ZKP development. Additionally, we enlist and discuss the currently deployed and in-deployment application software that fall under the different blockchain and non-blockchain categories. The second survey by Sun et al. [13] takes a narrower approach, concentrating on the intersection of ZKPs with blockchain technology. It outlines the security challenges within blockchain’s public and transparent nature and discusses how ZKPs can mitigate risks of private data exposure. Our work differentiates itself from these prior surveys by providing a more comprehensive practical application-oriented overview that covers both blockchain and non-blockchain applications of ZKPs in a broader context complemented by extensive analysis of the ZKP infrastructure being developed, including zkVMs, DSLs, and supportive technologies to guide developers to essential tools for building ZK-enabled applications efficiently.
This survey article serves as a comprehensive guide for practitioners and researchers, focusing on the broad range of practical applications and use cases of ZKPs. It systematically highlights the best practices, potential challenges, and the underlying impact of ZKPs on diverse application fields. Practitioners will find it a valuable roadmap that offers insights into incorporating ZKPs into various sophisticated digital systems. For researchers, it provides an overview of the current state of ZKP deployment, aiming to catalyze further research and development. By bridging the gap between theoretical cryptographic knowledge and practical application domains of ZKPs, this survey intends to enable professionals across disciplines to leverage the nuanced benefits of ZKPs in their work.
Figure 1: Survey Structure
@@
We organize the remainder of this article to methodically illustrate the intricacies of zkSNARKs and their wide range of applications. Section II characterizes the foundational components and requisite properties of SNARKs. Section III describes the infrastructure underpinning Zero-Knowledge Proofs, encompassing virtual machines, domain-specific languages, and pertinent libraries. Next, Section IV provides an extensive dissection of blockchain applications, which reviews the role of ZKPs across various blockchain layers and functionalities. Section V extends the discussion to non-blockchain application domains, detailing ZKPs’ utility in diverse contexts such as machine learning and digital identity verification. Finally, Section VI puts forward the future directions and discusses the expanding scope of ZKPs, then wraps up with a summary of our discussion and gives a reflective analysis of the presented content.
II Components of zkSNARKs
@@
Zero-Knowledge Proofs (ZKPs) are a cryptographic tool that enables one party, known as the prover, to prove to another party, the verifier, that a certain statement is true without revealing any information beyond the validity of the statement itself [14]. A “statement” in the context of ZKPs refers to a claim that can be mathematically verified, such as “I know two factors to a large number” or “I have executed a set of instructions from a virtual machine (VM) instruction set correctly”. This introduction aims to clarify the relationship between the broader category ZKPs and a particular subset known as Succinct Non-interactive Arguments of Knowledge (SNARKs), which aims to simplify the verification process by eliminating the need for interaction between two parties [^1] Our introduction is necessarily brief, for a more thorough discussion, we refer the reader to more detailed explainers on zk-SNARKs such as [11]..
There are several useful abstractions with which one could view the components that enable SNARKs. We will explore two such abstractions. First, we will follow the life-cycle of compiling high-level code into a provable form. Once the computational representation becomes provable, we will work with an interactive protocol, a process where two parties, referred to as a prover and verifier, exchange a series of messages to form a SNARK. First, we will explain the properties that a SNARK must adhere to.
II-A Definition and Properties of SNARKs
SNARKs, an acronym that stands for Succinct Non-interactive Arguments of Knowledge, represent a subset of zero-knowledge proofs with unique characteristics. Defining the acronym can help in highlighting these attributes:
  • Succinctness: SNARKs are distinguished by their compact proof size, which remains small regardless of the computational complexity or the size of the input data. This property is instrumental in environments where bandwidth or storage is limited, such as blockchain networks, ensuring that proofs can be transmitted and verified efficiently.
  • Non-interactivity: Unlike some Zero-Knowledge Proofs that require back-and-forth communication between the prover and the verifier, SNARKs are non-interactive. This means that the prover can generate a single proof that the verifier can independently check without further communication. Non-interactivity is achieved through the use of a common reference string shared between the prover and verifier.
  • Arguments of Knowledge: SNARKs assure not only that a certain statement is true but also that the prover possesses explicit knowledge of the information substantiating that statement. This aspect of SNARKs ensures that the proof carries with it a guarantee of knowledge, which is crucial for applications that require authentication or verification of authority.
SNARKs have several defining properties that ensure their functionality and security. These properties include:
@@ II-A Definition and Properties of SNARKs
    Completeness: If a statement is true and both the prover and verifier follow the protocol honestly, the verifier will be convinced of this truth by a valid proof. Mathematically, for every valid deterministic input $x$ and corresponding non-deterministic input $w$, often referred to as a witness, there exists a proof $\pi$ such that the verifier accepts $\pi$ as valid.
@@
    Soundness: If a statement is false, no cheating prover can convince the verifier of its truth. This is often a computational property, meaning that no efficient (polynomial time) prover can create a proof for a false statement that the verifier would accept.
@@
    Knowledge Soundness: Ensures that if a prover can convince a verifier of the truth of a statement, then the prover actually knows a private input for that statement. This distinguishes SNARKs from other non-interactive proof systems where a prover might be able to convince a verifier without actually knowing the private input.
@@
    Zero Knowledge: The proof does not reveal any information about the witness or the statement beyond its validity. There are different levels of zero-knowledge:
    • Computational Zero Knowledge: The verifier learns nothing from the proof that they couldn’t compute on their own.
    • Statistical Zero Knowledge: The verifier learns almost nothing from the proof, except with a small statistical difference.
    • Perfect Zero Knowledge: The proof gives absolutely no information to the verifier about the witness or statement, beyond the fact that the statement is true.
It’s worth noting that the differences between computational, statistical, and perfect properties revolve around the nature of the information leakage (or lack thereof) and the strength of the security guarantee.
II-B Lifecycle of a SNARK: From Python to Polynomials
II-B1 Frontends: From High-level code to circuits
This section explores the transition from high-level programming constructs to the more abstract representation as circuits, which are instrumental to the construction of a SNARK.
Consider the following Python function [^2] A Practical Caveat: While the transition from Python code to an arithmetic circuit illustrates the conceptual process underlying the construction of SNARKs, it’s important to note that, in practice, directly translating high-level programming languages to circuit representations may not be computationally efficient. To address this, developers often utilize zero-knowledge domain-specific languages (often referred to as zk-DSLs) or similar frameworks specifically designed for creating zero-knowledge proofs. These will be explored later in section III-B.:
def function(a, b, c):
    square_a = a * a
    square_b = b * b
    square_c = c * c
    multiply_a2b2 = square_a * square_b
    multiply_b2c2 = square_b * square_c
    sum_terms = multiply_a2b2 + multiply_b2c2
    return sum_terms
This function calculates the evaluation of a three-variate polynomial $f(a,b,c)=a^{2}b^{2}+b^{2}c^{2}$. Next, we represent this function as an arithmetic circuit. Arithmetic circuits are structured mathematical representations that decompose more-complex computations (such as the Python function above) into simple arithmetic operations (like addition and multiplication). This transformation, when coupled with the later transformations described below, will enable the properties of privacy and succinctness that enable the practicality of SNARKs.
[…]
Figure 2: An arithmetic circuit representation
In the context of cryptographic circuits like the one in Figure 2, the elements $a$, $b$, and $c$ are typically considered to be elements of a finite field $\mathbb{F}_{p}$, specifically numbers modulo a large prime number p. This use of a finite field ensures that all operations are performed within a closed, discrete set of numbers in the set {0, …, p-1}, providing properties that help maintain cryptographic security.
II-B2 Arithmetization: From circuits to matrices

### v2

@@ Abstract
Zero-knowledge proofs (ZKPs) enable computational integrity and privacy by allowing one party to prove the truth of a statement without revealing underlying data. Compared with alternatives such as homomorphic encryption and secure multiparty computation, ZKPs offer distinct advantages in universality and minimal trust assumptions, with applications spanning blockchain systems and confidential verification of computational tasks. This survey provides a technical overview of ZKPs with a focus on an increasingly relevant subset called zkSNARKs. Unlike prior surveys emphasizing algorithmic and theoretical aspects, we take a broader view of practical deployments and recent use cases across multiple domains including blockchain privacy, scaling, storage, and interoperability, as well as non-blockchain applications such as voting, authentication, timelocks, and machine learning. To support consistent comparison, we provide (i) a taxonomy of application areas, (ii) evaluation criteria including proof size, prover and verifier time, memory, and setup assumptions, and (iii) comparative tables summarizing key tradeoffs and representative systems. The survey also covers supporting infrastructure, including zero-knowledge virtual machines, domain-specific languages, libraries, and frameworks. While emphasizing zkSNARKs for their prevalence in deployed systems, we compare them with zkSTARKs and Bulletproofs to clarify transparency and performance tradeoffs. We conclude with future research and application directions.
Keywords zero-knowledge proofs $\cdot$ zkSNARKs $\cdot$ zkSTARKs $\cdot$ bulletproofs $\cdot$ blockchain $\cdot$ cryptography $\cdot$ privacy
1 Introduction
@@
Zero-knowledge proofs (ZKPs) are a set of cryptographic methods allowing one party to prove the validity of a claim to another without disclosing any of the claim’s underlying details. The seminal work by Goldwasser, Micali, and Rackoff [38] laid the groundwork for ZKPs in the 1980s. It introduced the principle of knowledge complexity as a metric for quantifying the information transferred from the prover to the verifier. Subsequently, Goldreich and others [36, 10] published several works until the 1990s that expanded the scope of ZKPs to a broader range of computational problems. They demonstrated their applicability to NP-complete problems under specific cryptographic assumptions. Another significant advancement in the development of ZKPs came with the introduction of succinctness in [14, 40] following the design paradigm of Kilian’s seminal paper [48]. The succinct and non-interactive properties of zkSNARKs greatly enhance their practical applicability in blockchain and non-blockchain systems. These properties make ZKPs applicable to digital systems that require both security and privacy [108]. This survey is written for three primary audiences: (i) Systems and security researchers can use it to understand the application landscape and recurring engineering bottlenecks that shape real deployments, (ii) Blockchain engineers can use it to navigate how ZKPs change system design choices around scalability, privacy, and cost, (iii) Students and applied cryptography practitioners can use it as a guided map from core components to modern toolchains and application patterns.
@@
The advent of ZKPs has represented a significant leap forward in the convergence of privacy and verifiability, which has attracted significant interest from both the cryptographic community and industry. It introduced a unique approach in the field of cryptography, differentiating itself from other privacy-preserving computation methods for distributed systems, such as homomorphic encryption and secure multiparty computation. While these methods are also being actively developed and advanced, each serves a specific purpose in information verification and privacy preservation. Homomorphic encryption enables computations on encrypted data without needing to decrypt, thereby preserving confidentiality while allowing the derivation of useful insights [85]. Secure multiparty computation enables trustless collaborations and allows parties to jointly compute a function over their inputs while keeping those inputs private [106]. ZKPs offer a distinct tradeoff profile, often enabling succinct public verification and privacy under assumptions that vary by proof family, by proving the truth of a statement without disclosing any other information.
@@
In modern digital systems, a tradeoff exists between openness and privacy. Blockchains, for example, prioritize transparency to ensure trust and prevent fraud, with every transaction openly verifiable. However, this transparency can compromise privacy [71]. Despite the pseudonymous nature of blockchain transactions, advanced analytics can de-anonymize users by correlating on-chain and off-chain data. Such exposure can reveal a user’s entire transaction history, leading to privacy breaches and potential targeted threats. As digital infrastructures become increasingly complex, striking the right balance between transparency for security and preserving user privacy becomes a vital research challenge. In this context, ZKPs emerge as a robust solution addressing the challenges posed by such tradeoffs in digital systems.
@@
ZKPs enable users to verify private data such as bank balances or credit scores without revealing specifics. Furthermore, they can ensure anonymity in authorization processes by allowing access to restricted areas or proving regional origin without disclosing detailed credentials. In the financial domain, ZKPs can allow for identity-free payments and tax submissions without revealing exact earnings. Additionally, they facilitate trustless outsourcing by letting organizations validate results without redoing the entire computation. Furthermore, ZKPs can modify blockchain operations by shifting computation off-chain while preserving network-wide verification. Overall, ZKPs exemplify the merging of verification with privacy across diverse applications [76].
@@
The practical applications of ZKPs rest on verifiable computation: the ability to prove that a computation was performed correctly without revealing the inputs or the computation process. Verifiable computation enables the two main value propositions of ZKPs – succinctness and privacy. In the remainder of this survey, we will see how each of these two values contributes to the practical applications of ZKPs. Succinctness in ZKPs allows for the quick verification of the correctness of a computation without the extensive resources typically required for direct computation execution. This matters especially in contexts where computational resources are limited or expensive, such as blockchain networks. For instance, ZKPs enable the validation of transaction blocks or smart contract executions without necessitating each node to replay the entire sequence of transactions or computations. This aspect dramatically reduces the computational load on the network, enhancing scalability and throughput. Privacy, the second value proposition, follows from the ability of ZKPs to prove correctness without revealing the underlying data. This property is especially useful when sensitive or confidential data is involved, enabling applications such as private voting systems and confidential financial transactions in which process integrity is maintained without exposing participants’ information.
While zkSNARKs are emphasized because they are widely used in many deployed systems, especially where small proofs and cheap verification are important, this survey covers ZK applications broadly and compares major proof families, including zkSNARKs, zkSTARKs, and Bulletproofs, to explain when different assumptions and performance profiles are preferred in practice. Throughout the survey, we use a consistent evaluation lens to compare systems and application designs, focusing on proof size, prover time, verifier time, peak memory, setup assumptions (trusted versus transparent), and, for blockchain deployments, on-chain verification cost and data availability cost.
@@
In the ZKP literature, there exist two key survey articles relevant to our work. The first survey by Morais et al. [65] dives deep into Zero-Knowledge Range Proofs (ZKRP) with a particular emphasis on the algorithmic details required for the implementation of ZKRPs. Furthermore, it defines the fundamental theoretical background behind ZKRPs and describes possible generic and blockchain-specific use cases of ZKRPs. In contrast, our survey covers a broader range of application categories and use cases. Further, we provide an overview of the underlying theoretical components, tools, and infrastructure needed for ZKP development. Additionally, we enlist and discuss currently deployed and in-deployment application software that fall under different blockchain and non-blockchain categories. The second survey by Sun et al. [100] takes a narrower approach, concentrating on the intersection of ZKPs with blockchain technology. It outlines the security challenges within blockchain’s public and transparent nature and discusses how ZKPs can mitigate risks of private data exposure. Our work differentiates itself from these prior surveys by providing a more comprehensive practical application-oriented overview that covers both blockchain and non-blockchain applications of ZKPs, complemented by an analysis of the ZKP infrastructure being developed, including zkVMs, DSLs, and supportive technologies to guide developers to essential tools for building ZK-enabled applications efficiently.
This survey provides a comprehensive overview for practitioners and researchers, covering the practical applications and use cases of ZKPs. It describes best practices, challenges, and the impact of ZKPs across application fields. For practitioners, it offers guidance on incorporating ZKPs into real systems. For researchers, it summarizes the current state of ZKP deployment and identifies open problems. By connecting the theoretical cryptographic foundations to practical application domains, the survey aims to help professionals across disciplines apply ZKPs effectively.
@@
We organize the remainder of this article to methodically illustrate the components, toolchains, and applications of ZKPs using a consistent evaluation lens. Section 2 defines core primitives, terminology, and the foundational components and requisite properties used throughout the survey. Section 3 describes the infrastructure underpinning ZKPs, encompassing virtual machines, domain-specific languages, and pertinent libraries. Next, Section 4 reviews blockchain applications and compares system designs using the evaluation criteria described above. Section 5 extends the discussion to non-blockchain application domains, detailing ZKPs’ utility in contexts such as machine learning and digital identity verification. Finally, Section 6 discusses open challenges and future directions, then wraps up with a summary and reflective analysis. The overall survey structure is depicted in Figure 1.
ZK Developer ToolsZK VirtualMachinesZK Domain-SpecificLanguagesZK Libraries &FrameworksZK HardwareAccelerationZKP ApplicationsBlockchain ApplicationsNon-Blockchain ApplicationsProof ofIdentityLayer 1BlockchainsSupply Chain /EnterpriseLayer 2ScalingInteroperabilityTransactionPrivacyStorageProof ofReservesProof of IdentityZK ML/AIOther ApplicationsImplemented using
Figure 1: ZKP Applications & Developer Tools (Survey Structure).
2 Foundations of Zero-Knowledge Proofs
@@
Zero-knowledge proofs (ZKPs) are cryptographic protocols that enable one party, known as the prover, to convince another party, the verifier, that a statement is true without revealing any information beyond the validity of the statement itself [38]. In this context, a statement is a claim that can be efficiently verified, such as knowledge of factors of a large integer, or correct execution of a program under a specified instruction set. This section clarifies the relationship between the broad class of ZKPs and two widely used refinements: SNARKs and zkSNARKs  [^1] Our overview is necessarily brief. For a more detailed discussion, we refer the reader to explainers such as [76]..
2.1 Definitions and Core Properties
To avoid ambiguity, we use the following layered terminology throughout the survey.
  • ZKP: A zero-knowledge proof is a protocol for proving that a statement is true without revealing the witness. ZKPs may be interactive or non-interactive, and they may be succinct or non-succinct depending on the construction.
  • SNARK: A SNARK is a succinct and non-interactive argument of knowledge. Importantly, being a SNARK does not automatically imply the zero-knowledge property. A SNARK may or may not be instantiated with a zero-knowledge guarantee, depending on how it is constructed.
  • zkSNARK: A zkSNARK is a SNARK that additionally satisfies the zero-knowledge property, meaning that the proof reveals nothing about the witness beyond the fact that the statement is true.
The term argument distinguishes computational soundness from information-theoretic soundness. In a proof system with information-theoretic soundness, no prover (even unbounded) can convince the verifier of a false statement. In an argument system, soundness is computational: it holds against any efficient (polynomial-time) prover under standard cryptographic assumptions. SNARKs are arguments because their security relies on computational assumptions.
We summarize the most commonly referenced properties below. When a construction is specifically a zkSNARK, it includes the zero-knowledge property in addition to succinctness, non-interactivity, and knowledge soundness.
@@ 2.1 Definitions and Core Properties
    Completeness: If the statement is true and both parties follow the protocol, the verifier accepts a valid proof.
@@
    Soundness (computational): If the statement is false, no efficient prover can convince the verifier to accept, except with negligible probability.
@@
    Knowledge soundness: If a prover convinces the verifier, then the prover knows a valid witness for the statement (formalized via an extractor).
@@
    Zero knowledge (when present): The proof leaks no information about the witness beyond statement validity. Depending on the construction, this can be computational, statistical, or perfect.
Many deployed zkSNARK systems rely on public parameters generated in a setup phase. A common reference string (CRS) refers to these public parameters that are shared by prover and verifier. In some constructions, the CRS is generated together with secret randomness (often called a trapdoor). If the trapdoor is ever learned by an adversary, it may enable forging proofs or breaking soundness, depending on the specific scheme. This is why CRS generation is often done via a multi-party ceremony designed so that as long as one participant is honest and discards their secret randomness, the trapdoor remains unknown. In contrast, transparent setups aim to avoid secret trapdoors by deriving public parameters from public randomness and standard assumptions, so that the setup does not rely on the secrecy of any toxic waste. Ongoing standardization efforts, such as those coordinated by ZKProof.org [111], are formalizing setup categories and interoperability requirements across proof systems.
Fiat–Shamir versus setup: Non-interactivity and setup are related but distinct concepts. The Fiat–Shamir heuristic [30] is a transform that turns certain interactive proof protocols into non-interactive ones by replacing the verifier’s random challenges with outputs of a hash function (typically modeled as a random oracle). This addresses interactivity. Separately, trusted setup versus transparent setup concerns how public parameters are generated and whether they depend on secret trapdoor material. A system can be non-interactive while still requiring a trusted setup, and it can also be non-interactive with a transparent setup, depending on the proof family.
Table 1: High-level tendencies of common proof-system families used in practice. These entries are broad tendencies rather than universal properties.
Category Zero Knowledge Succinct Interactive Setup Post-quantum
ZKP (general) yes optional optional optional depends
SNARK optional yes no trusted no
zkSNARK yes yes no trusted no
zkSTARK yes yes no transparent yes
Bulletproofs yes no no transparent typically no
2.2 Lifecycle of a SNARK: From Python to Polynomials
Figure 2: ZKP Pipeline: The frontend compiles programs into a constraint system; the backend combines a polynomial IOP with a polynomial commitment scheme (PCS) producing the succinct proof and verification key.
2.2.1 Frontends: From High-level code to circuits
The Python function [^2] A practical caveat: While the transition from Python code to an arithmetic circuit illustrates the conceptual process underlying SNARK construction, directly translating a general-purpose language into constraints is often inefficient. In practice, developers commonly use zero-knowledge domain-specific languages (zkDSLs) and compilers designed for constraint generation. These are explored in Section 3.2. (Figure 2) calculates the evaluation of a three-variate polynomial $f(a,b,c)=a^{2}b^{2}+b^{2}c^{2}$. Next, we represent this function as an arithmetic circuit. Arithmetic circuits decompose complex computations into simple arithmetic operations (such as addition and multiplication). This transformation, together with later steps described below, enables the privacy and succinctness properties that make modern SNARK systems practical.
In cryptographic circuits, the elements $a$, $b$, and $c$ are typically treated as elements of a finite field $\mathds{F}_{p}$, that is, integers modulo a large prime $p$. This ensures that all operations are performed within a closed set $\{0,\ldots,p-1\}$, enabling algebraic techniques used by modern proving systems.
2.2.2 Arithmetization: From circuits to matrices
Table 2: High-level comparison of arithmetization strategies.
Property R1CS Plonkish AIR CCS (folding)
Constraint style Bilinear constraints on linear combos Custom gates + permutation/copy constraints Polynomial constraints over an execution trace Generalized constraints designed for folding/IVC
Strengths Simple, mature tooling Flexible arith., efficient lookups Transparent (no trusted setup), good for large traces Enables efficient folding / incremental verification
Typical prover cost driver MSMs (and pairings in some systems) FFT/NTT + MSM Hashing + large-trace memory/IO Backend-dependent
Lookups Usually emulated (or via extra gadgets) Native (Plookup/LogUp/Lasso, etc.) Via lookup arguments (e.g., LogUp) Via composition within the folded system
Recursion / IVC Often pairing-cycle-based in pairing SNARKs Accumulation / recursion-friendly designs FRI-style recursion/composition Folding (Nova, HyperNova, etc.)
Examples Groth16, Marlin Plonk, Halo2 STARKs, Cairo Nova, HyperNova